2 3 4 5 6 8 9 10 Solve the right triangle. Round your answers to the nearest tenth. B = U c=0 X 5 ? Continue
Answer #1

Given 41A 6-17 an 6:17 LA = 41 A ABC LA+CB+22=180 41+ LB +90=180 LB : 180-131 LB-49 Ang tan A 96 tan 41 = 14.8 a =17 ton 41 =